OpenStudy (anonymous):

find the modulus and argument of the complex number 5-i / 2-3i

OpenStudy (anonymous):

Firstly rationalize the denominator... Multiply numerator and denominator with conjugate of 2 - 3i ...

OpenStudy (anonymous):

okay .. thanks :)

OpenStudy (shubhamsrg):

you can do that but there is a simpler method also : for z= p/q where z,p,q are all complex nos, |z| = | p/q | = |p| / |q| and arg(z) = arg( p/q ) = arg(p) - arg(q) both shall yield the same ans